OpenStudy (anonymous):

The total U.S. consumption of electricity in a particular year was 5.0 x 1012 kilowatt-hours. What is the mass equivalence of this amount of energy? Show all calculations leading to an answer.

OpenStudy (anonymous):

5.0x10^12

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

Honestly, I've never seen this before, but I think you convert kW-h to J and use E = mc²

OpenStudy (anonymous):

quite right ....just calculate the year's energy in joules and then use E=mc^2 to find out the mass equivalence ...

OpenStudy (anonymous):

can you walk me through changing it to joules please

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

1 kilowatt hour = 3 600 000 joules

OpenStudy (anonymous):

convert it your Kwh into joules then put it in E=MC^2

OpenStudy (anonymous):

E being your energy in jouls ....and C being the speed of light

OpenStudy (anonymous):

use SI units

OpenStudy (anonymous):

so 5.0*10^12=5,000,000,000,000

OpenStudy (anonymous):

now multiply it by 3 600 000

OpenStudy (anonymous):

yu get your enerygy in jouls then

OpenStudy (anonymous):

don t get rid of the exponantional notation here ...it help keeps the number small

OpenStudy (anonymous):

so 1.8*10^19j

OpenStudy (anonymous):

3 600 000 x (your enrgy in Kwh) = MC^2 you need to fing out M

OpenStudy (anonymous):

J*

OpenStudy (anonymous):

yeah ... there ya go ^_^

OpenStudy (anonymous):

and than the equation i need is M=E/C^2

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good work

OpenStudy (anonymous):

So M=1.8*10^19J/299792458m/s

OpenStudy (anonymous):

gottta change m/s to J

OpenStudy (anonymous):

Basically, the equation states that matter and energy are identical. If you could convert a small amount of matter completely, it would yield a tremendous amount of energy. in your case how ever you are finding the mass cause you already know the energy

OpenStudy (anonymous):

everything should be in SI units ....like meter for distance jouls for energy ...

OpenStudy (anonymous):

so what do i do than

OpenStudy (anonymous):

So M=1.8*10^19J/299792458m/s you said this was your answer right ?

OpenStudy (anonymous):

yes but isnt there something more i have to do

OpenStudy (anonymous):

that figure is your answer .... express it in kilograms and your all set

OpenStudy (anonymous):

how do i express it in kilograms

OpenStudy (anonymous):

just do 1.8*10^19/299792458 and add the alphabets Kg in the end

OpenStudy (anonymous):

would you believe it i am driving right now ....and i pull over to the side to answer your questions

OpenStudy (anonymous):

are you clear on this ?

OpenStudy (anonymous):

haha thank you so much

OpenStudy (anonymous):

feel free to ask .... i am in the parking lot now

OpenStudy (anonymous):

Would you help me with one more question if i post it

OpenStudy (anonymous):

No problem bro

OpenStudy (anonymous):

depends if i can solve it myself .... ... hit me with it